OpenStudy (horrorfreakk):

PLEASE HELP ME!! I WILL MEDAL. <3 I need help doing this assignment on FLVS. 03.03 Linear Functions. The instructions will be in the cooments since there's no room on the status.

OpenStudy (horrorfreakk):

Project Option 1—Individually Sal's Sandwich Shop sells wraps and sandwiches as part of its lunch specials. The profit on every sandwich is $2 and the profit on every wrap is $3. Sal made a profit of $1,470 from lunch specials last month. The equation 2x + 3y = 1,470 represents Sal's profits last month, where x is the number of sandwich lunch specials sold and y is the number of wrap lunch specials sold. Change the equation to slope-intercept form. Identify the slope and y-intercept of the equation. Be sure to show all your work. Describe how you would graph this line using the slope-intercept method. Be sure to write using complete sentences. Write the equation in function notation. Explain what the graph of the function represents. Be sure to use complete sentences. Graph the function. On the graph, make sure to label the intercepts. You may graph your equation by hand on a piece of paper and scan your work or you may use graphing technology. Suppose Sal's total profit on lunch specials for the next month is $1,593. The profit amounts are the same: $2 for each sandwich and $3 for each wrap. In a paragraph of at least three complete sentences, explain how the graphs of the functions for the two months are similar and how they are different. Below is a graph that represents the total profits for a third month. Write the equation of the line that represents this graph. Show your work or explain how you determined the equations. graph of line going through ordered pairs (0, 300) and (450, 0) 03.03 Linear Functions—Option 1 Rubric Requirements Possible Points Student Points Student changes equation to slope-intercept form. Student shows all work and identifies the slope and y-intercept of the equation. 4 Student writes a description, which is clear, precise, and correct, of how to graph the line using the slope-intercept method. 4 Student changes equation to function notation. Student explains clearly what the graph of the equation represents. 2 Student graphs the equation and labels the intercepts correctly. 4 Student writes at least three sentences explaining how the graphs of the two equations are the same and how they are different. 4 Student writes the equation of the graph in either form correctly and shows work or explains the process. 2 Print

OpenStudy (retireed):

Change the equation to slope-intercept form. Identify the slope and y-intercept of the equation. Be sure to show all your work. 2x + 3y = 1,470 in the y-intercept form y = mx + b Convert the equation to y-intercept form

OpenStudy (horrorfreakk):

That's the answer to the full problem?